If the Intensity of the Sun Light hitting the Earth is 1,399 W/m2, What is the Magnitude of the Electric field?
Answer #1

Given that -7 8 m/s Mo = 12.56X10 Time A C- 3x10 / I = 1399 W/m2 2 E I Glo (3xing) (12.56x137)1299) o I =) Mo E- N/ E = 726